A plumer gave an estimate for the renovation of a kitchen. Her hourly pay is $27 per hour and the plumbing parts will cost $80. If her total estimate is $404, how many hours does she expect this job to take?
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

404-80=324/27=12 hrs

Answer by checkley77(12844)

404-80=27x
324=27x
x=324/27
x=12 hours.